# Classification of thin Jordan schemes

### v2

@@ 4. Coherent closure of thin Jordan schemes
Let $\mathfrak{X}=(\Omega,C)$ be the coherent closure of $G$. Since $H$ centralizes $G$, each $h\in H$ is an automorphism of every $g\in G$. Therefore every basic relation of $\mathfrak{X}$ is a union of $2$-orbits of $H$. We must show each $2$-orbit of $H$ is a union of basic relations of $\mathfrak{X}$.
@@
The relation $r_{\omega_{1},\omega_{2}}:=\bigcap_{g\in G_{\omega_{1}\rightarrow\omega_{2}}}g$ is nonempty and is the union of certain elements from $C$. We claim it is a $2$-orbit of $H$. Fix a permutation $g\in G_{\omega_{1}\rightarrow\omega_{2}}$ and write $r_{\omega_{1},\omega_{2}}=gr_{\omega_{1},\omega_{1}}$. Obviously $r_{\omega_{1},\omega_{1}}=1_{\Omega_{1}}$ where $\Omega_{1}=\mathsf{Fix}(G_{\omega_{1}})$. Clearly $\Omega_{1}$ is $H$-invariant implying that $1_{\Omega_{1}}$ is $H$-invariant as well.
@@
Finally, it is well known that $|\Omega_{1}|=[N_{G}(G_{\omega_{1}}):G_{\omega_{1}}]=|C_{{\sf Sym}(\Omega)}(G)|=|H|$. Hence $H$ acts transitively in $1_{\Omega_{1}}$, and therefore it also acts transitively in $g1_{\Omega_{1}}=r_{\omega_{1},\omega_{2}}$. We conclude that $r_{\omega_{1},\omega_{2}}$ is a $2$ orbit of $H$. ∎
@@
Let $(S,\diamond)$ be a loop, and let $G\subseteq{\sf Sym}(S)$ be the subgroup generated by $\{\ell_{s}\}_{s\in S}$. Then $C_{{\sf Sym}(S)}(G)$ coincides with the subgroup $\widehat{N}_{\rho}:=\{r_{s}\}_{s\in N_{\rho}}$ of ${\sf Sym}(S)$, where $N_{\rho}$ is the right nucleus of $(S,\diamond)$.
@@ 4.1. Jordan algebraic automorphisms of coherent configurations
Proof of (a). Let $\Omega_{1},\dots,\Omega_{k}$ be a complete set of $H$-orbits on $\Omega$. Clearly $|\Omega_{i}|=|H|$ for all $i\in[k]$. Note that $\Omega_{1}\times[k]$ is the point set of $(\Omega_{1},C\mathord{\upharpoonright}_{\Omega_{1}})\otimes\mathfrak{D}_{k}$. Thus, to build a bijection $f:\Omega_{1}\times[k]\rightarrow\Omega$ we choose arbitrary $\omega_{i}\in\Omega_{i},i\in[k]$. Define $f(\omega,i)=h(\omega_{i})$ where $h\in H$ is the unique element satisfying $\omega=h(\omega_{1})$.
We now check that $f$ maps basic relations of $(\Omega_{1},C\mathord{\upharpoonright}_{\Omega_{1}})\otimes\mathfrak{D}_{k}$ onto basic relations of $\mathfrak{C}$. Since in both configurations all basic relations have the same cardinality, namely $|H|$, it is sufficient to prove that for any quadruple of points $(\alpha,i),(\beta,j),(\alpha^{\prime},i^{\prime}),(\beta^{\prime},j^{\prime})\in\Omega_{1}\times[k]$ the following holds:
@@
It follows from the first equality that $i^{\prime}=i$ and $j^{\prime}=j$. Also $\alpha^{\prime}=h(\alpha)$ and $\beta^{\prime}=h(\beta)$ where $h$ is the unique element of $H$ that maps $\alpha$ to $\alpha^{\prime}$. Let $a,b\in H$ be elements satisfying $\alpha=a(\omega_{1}),\beta=b(\omega_{1})$. Then $\alpha^{\prime}=ha(\omega_{1})$ and $\beta^{\prime}=hb(\omega_{1})$. Then we obtain $f(\alpha,i)=a(\omega_{i})$, $f(\beta,j)=b(\omega_{j})$ and $f(\alpha^{\prime},i)=ha(\omega_{i})$, $f(\beta^{\prime},j)=hb(\omega_{j})$. The latter equality implies $(f(\alpha^{\prime},i),f(\beta^{\prime},j))=(h(f(\alpha,i)),h(f(\beta,j)))$, whence we have $r(f(\alpha,i),f(\beta,j))=r(f(\alpha^{\prime},i^{\prime}),f(\beta^{\prime},j^{\prime}))$ proving (a).
Proof of (b). We first describe the group of isomorphisms $\mathsf{Iso}(\mathfrak{C})$ (referred to as color automorphisms in [10]). Since $H$ acts semiregularly on $\Omega$, $H$ is $2$-closed [18]. In this case $\mathsf{Iso}(\mathfrak{C})$ coincides with the normalizer of $H$ in the symmetric group ${\sf Sym}(\Omega)$.
@@
$E_{\ell,j}=2E_{i,j}\star E_{\ell,i}=2\phi(E_{j,i})\star\phi(E_{\ell,i})=2\phi(E_{j,i}\star E_{\ell,i})=0,$
@@
$E_{i,j}=2E_{i,\ell}\star E_{\ell,j}=2\phi(E_{\ell,i})\star\phi(E_{\ell,j})=2\phi(E_{\ell,i}\star E_{\ell,j})=0.$
@@
It follows from above that $\phi$ acts on ${\mathcal{A}}_{0}\cong M_{k}(\mathbb{C})$ either trivially or via matrix transposition. Combining $\phi$ with t in the latter case, we may assume $\phi$ acts trivially on ${\mathcal{A}}_{0}$, i.e. $\phi(E_{i,j})=E_{i,j}$ for all $i,j\in[k]$. This implies the set of relations $\{[f]_{i,j}\}_{f\in F}$ is $\phi$-invariant for any $i,j\in[k]$, i.e. there exists a permutation $\phi_{ij}\in{\sf Sym}(F)$ such that $\phi([f]_{i,j})=[\phi_{ij}(f)]_{i,j}$.
It follows from $f]_{ii}\star[h]_{ii}=[fh]_{ii}+[hf]_{ii}$ that $\phi_{ii}(fh)\in\{\phi_{ii}(f)\phi_{ii}(h),\phi_{ii}(h)\phi_{ii}(f)\}$. Thus $\phi_{ii}$ is a half-automorphism [15] of $F$. According to [15], $\phi_{ii}$ is either an automorphism or an anti-automorphism.
@@ 4.2. Thin regular Jordan schemes are autonomous
By way of contradiction, assume $\mathfrak{S}$ is the algebraic fusion of a coherent configuration $\mathfrak{C}=(S,C)$ with respect to a certain subgroup $\Phi\leq\mathsf{JAut}(\mathfrak{C})$. Then $\hat{s}=\Phi(c)$ for every $s\in S$ and $c\in C$ contained in $\hat{s}$. In particular, this implies that each $c\in C$ is thin. By [5, Exercise 2.7.35], every such configuration is Schurian, i.e. there exists $H\leq{\sf Sym}(S)$ such that $C$ is the set of $2$-orbits of $H$. Since every $2$-orbit of $H$ is thin, we have that $H$ is semiregular. Thus by Proposition 4.3, $\mathsf{JAut}(\mathfrak{C})=\mathsf{TAut}(\mathfrak{C})$. Hence the subgroup $\Phi_{0}:=\Phi\cap\mathsf{AAut}(\mathfrak{C})$ has index at most two in $\Phi$. Moreover, since $\Phi_{0}\leq\mathsf{AAut}(\mathfrak{C})$, the rainbow $^{\Phi_{0}}\mathfrak{C}=(S,C_{0})$ where $C_{0}:=\{\Phi_{0}(c)\}_{c\in C}$ is a coherent configuration. Thus $\hat{S}={}^{\Phi}C\sqsubseteq C_{0}\sqsubseteq C$. Since $\mathfrak{S}$ is not a coherent configuration, the inclusion $\hat{S}\sqsubseteq C_{0}$ must be proper, and $\Phi:\Phi_{0}]=2$. In particular, $\Phi_{0}$ is normal in $\Phi$, which in turn implies that each $\Phi$-orbit on $C$ is either a $\Phi_{0}$-orbit or a disjoint union of two $\Phi_{0}$-orbits of equal size. Therefore, each $\hat{s}\in\hat{S}$ is either a basic relation of $^{\Phi_{0}}\mathfrak{C}$ or the union of two basic relations of $^{\Phi_{0}}\mathfrak{C}$ of equal size. Since $^{\Phi_{0}}\mathfrak{C}$ and $^{\Phi}\mathfrak{C}$ are distinct, there exists $\hat{s}\in\hat{S}$ which splits into a disjoint union of two basic relations of $^{\Phi_{0}}\mathfrak{C}$ of equal size, say $\hat{s}=c_{1}\cup c_{2}$. It follows from $|c_{1}|=|c_{2}|=|S|/2$ that both $c_{1}$ and $c_{2}$ are nonregular basic relations of $^{\Phi_{0}}\mathfrak{C}$. Therefore $^{\Phi_{0}}\mathfrak{C}$ is a nonhomogeneous coherent configuration, that is, each of its basic relations is nonregular. This in turn implies that all basic relations of $^{\Phi_{0}}\mathfrak{C}$ have size $|S|/2$.
@@
Acknowledgment
The authors wish to thank Misha Klin for helpful discussions and suggestions.

### v3

@@ 4. Coherent closure of thin Jordan schemes
Let $\mathfrak{X}=(\Omega,C)$ be the coherent closure of $G$. Since $H$ centralizes $G$, each $h\in H$ is an automorphism of every $g\in G$. Therefore every basic relation of $\mathfrak{X}$ is a union of $2$-orbits of $H$. We must show that each $2$-orbit of $H$ is a union of basic relations of $\mathfrak{X}$.
@@
The relation $r_{\omega_{1},\omega_{2}}:=\bigcap_{g\in G_{\omega_{1}\rightarrow\omega_{2}}}g$ is nonempty and is the union of certain elements from $C$. We claim $r_{\omega_{1},\omega_{2}}$ is a $2$-orbit of $H$. Fix a permutation $g\in G_{\omega_{1}\rightarrow\omega_{2}}$ and write $r_{\omega_{1},\omega_{2}}=gr_{\omega_{1},\omega_{1}}$. Obviously $r_{\omega_{1},\omega_{1}}=1_{\Omega_{1}}$ where $\Omega_{1}$ is the fixed-point set $\mathsf{Fix}(G_{\omega_{1}})$ of $\Omega$ under $G_{\omega_{1}}$. Clearly $\Omega_{1}$ is $H$-invariant implying that $1_{\Omega_{1}}$ is $H$-invariant as well.
@@
Finally, it is well known that $|\Omega_{1}|=[N_{G}(G_{\omega_{1}}):G_{\omega_{1}}]=|C_{{\sf Sym}(\Omega)}(G)|=|H|$. Hence $H$ acts transitively in $1_{\Omega_{1}}$, and therefore it also acts transitively in $g1_{\Omega_{1}}=r_{\omega_{1},\omega_{2}}$. We conclude that $r_{\omega_{1},\omega_{2}}$ is a $2$-orbit of $H$. ∎
@@
Let $(S,\diamond)$ be a loop, and let $G\leq Sym(S)$ be the subgroup generated by $\{\ell_{s}\}_{s\in S}$. Then $C_{{\sf Sym}(S)}(G)$ coincides with the subgroup $\widehat{N}_{\rho}:=\{r_{s}\}_{s\in N_{\rho}}$ of ${\sf Sym}(S)$, where $N_{\rho}$ is the right nucleus of $(S,\diamond)$.
@@ 4.1. Jordan algebraic automorphisms of coherent configurations
Proof of a: Let $\Omega_{1},\dots,\Omega_{k}$ be a complete set of $H$-orbits on $\Omega$. Clearly $|\Omega_{i}|=|H|$ for all $i\in[k]$. Note that $\Omega_{1}\times[k]$ is the point set of $(\Omega_{1},C\mathord{\upharpoonright}_{\Omega_{1}})\otimes\mathfrak{D}_{k}$. Thus, to build a bijection $f:\Omega_{1}\times[k]\rightarrow\Omega$ we choose arbitrary $\omega_{i}\in\Omega_{i},i\in[k]$. Define $f(\omega,i)=h(\omega_{i})$ where $h\in H$ is the unique element satisfying $\omega=h(\omega_{1})$.
We now check that $f$ maps basic relations of $(\Omega_{1},C\mathord{\upharpoonright}_{\Omega_{1}})\otimes\mathfrak{D}_{k}$ onto basic relations of $\mathfrak{C}$. Since in both configurations all basic relations have the same cardinality, namely $|H|$, it is sufficient to prove the following holds for any quadruple of points $(\alpha,i),(\beta,j),(\alpha^{\prime},i^{\prime}),(\beta^{\prime},j^{\prime})\in\Omega_{1}\times[k]$:
@@
It follows from the first equality that $i^{\prime}=i$ and $j^{\prime}=j$. Also $\alpha^{\prime}=h(\alpha)$ and $\beta^{\prime}=h(\beta)$ where $h$ is the unique element of $H$ that maps $\alpha$ to $\alpha^{\prime}$. Let $a,b\in H$ be elements satisfying $\alpha=a(\omega_{1}),\beta=b(\omega_{1})$ from which $\alpha^{\prime}=ha(\omega_{1})$ and $\beta^{\prime}=hb(\omega_{1})$ follows. Then we obtain $f(\alpha,i)=a(\omega_{i})$, $f(\beta,j)=b(\omega_{j})$ and $f(\alpha^{\prime},i)=ha(\omega_{i})$, $f(\beta^{\prime},j)=hb(\omega_{j})$. The latter equality implies $(f(\alpha^{\prime},i),f(\beta^{\prime},j))=(h(f(\alpha,i)),h(f(\beta,j)))$, whence we have $r(f(\alpha,i),f(\beta,j))=r(f(\alpha^{\prime},i^{\prime}),f(\beta^{\prime},j^{\prime}))$ proving (a).
Proof of b: We first describe the group of isomorphisms $\mathsf{Iso}(\mathfrak{C})$ (referred to as color automorphisms in [13]). Since $H$ acts semiregularly on $\Omega$, $H$ is $2$-closed [22]. In this case $\mathsf{Iso}(\mathfrak{C})$ coincides with the normalizer of $H$ in the symmetric group ${\sf Sym}(\Omega)$.
@@
$E_{\ell,j}=2E_{i,j}\star E_{\ell,i}=2\phi(E_{j,i})\star\phi(E_{\ell,i})=2\phi(E_{j,i}\star E_{\ell,i})=O,$
@@
$E_{i,j}=2E_{i,\ell}\star E_{\ell,j}=2\phi(E_{\ell,i})\star\phi(E_{\ell,j})=2\phi(E_{\ell,i}\star E_{\ell,j})=O.$
@@
It follows from above that $\phi$ acts on ${\mathcal{A}}_{0}\cong M_{k}(\mathbb{C})$ either trivially or via matrix transposition. Combining $\phi$ with $\tau$ in the latter case, we may assume $\phi$ acts trivially on ${\mathcal{A}}_{0}$, i.e. $\phi(E_{i,j})=E_{i,j}$ for all $i,j\in[k]$. This implies the set of relations $\{[f]_{i,j}\}_{f\in F}$ is $\phi$-invariant for any $i,j\in[k]$, i.e. there exists a permutation $\phi_{ij}\in{\sf Sym}(F)$ such that $\phi([f]_{i,j})=[\phi_{ij}(f)]_{i,j}$.
It follows from $2[f]_{ii}\star[h]_{ii}=[fh]_{ii}+[hf]_{ii}$ that $\phi_{ii}(fh)\in\{\phi_{ii}(f)\phi_{ii}(h),\phi_{ii}(h)\phi_{ii}(f)\}$. Thus $\phi_{ii}$ is a half-automorphism [19] of $F$. According to [19], $\phi_{ii}$ is either an automorphism or an anti-automorphism.
@@ 4.2. Thin regular Jordan schemes are autonomous
By way of contradiction, assume $\mathfrak{S}$ is the algebraic fusion of a coherent configuration $\mathfrak{C}=(S,C)$ with respect to a certain subgroup $\Phi\leq\mathsf{JAut}(\mathfrak{C})$. Then $\hat{s}=\Phi(c)$ for every $s\in S$ and $c\in C$ contained in $\hat{s}$. In particular, this implies that each $c\in C$ is thin. By [7, Exercise 2.7.35], every such configuration is Schurian, i.e. there exists $H\leq{\sf Sym}(S)$ such that $C$ is the set of $2$-orbits of $H$. Since every $2$-orbit of $H$ is thin, we have that $H$ is semiregular. Thus by Proposition 4.3, $\mathsf{JAut}(\mathfrak{C})=\mathsf{TAut}(\mathfrak{C})$. Hence the subgroup $\Phi_{0}:=\Phi\cap\mathsf{AAut}(\mathfrak{C})$ has index at most two in $\Phi$. Moreover, since $\Phi_{0}\leq\mathsf{AAut}(\mathfrak{C})$, the rainbow $^{\Phi_{0}}\mathfrak{C}=(S,C_{0})$ where $C_{0}:=\{\Phi_{0}(c)\}_{c\in C}$ is a coherent configuration. Thus $\hat{S}={}^{\Phi}C\sqsubseteq C_{0}\sqsubseteq C$. Since $\mathfrak{S}$ is not a coherent configuration, the inclusion $\hat{S}\sqsubseteq C_{0}$ must be proper, and $\Phi:\Phi_{0}]=2$. In particular, $\Phi_{0}$ is normal in $\Phi$, which in turn implies that each $\Phi$-orbit on $C$ is either a $\Phi_{0}$-orbit or a disjoint union of two $\Phi_{0}$-orbits of equal size. Therefore, each $\hat{s}\in\hat{S}$ is either a basic relation of $^{\Phi_{0}}\mathfrak{C}$ or the union of two basic relations of $^{\Phi_{0}}\mathfrak{C}$ of equal size. Since $^{\Phi_{0}}\mathfrak{C}$ and $^{\Phi}\mathfrak{C}$ are distinct, there exists $\hat{s}\in\hat{S}$ which splits into a disjoint union of two basic relations of $^{\Phi_{0}}\mathfrak{C}$ of equal size, say $\hat{s}=c_{1}\cup c_{2}$. It follows from $|c_{1}|=|c_{2}|=|S|/2$ that both $c_{1}$ and $c_{2}$ are nonregular basic relations of $^{\Phi_{0}}\mathfrak{C}$. Therefore $^{\Phi_{0}}\mathfrak{C}$ is a nonhomogeneous coherent configuration, that is, each of its basic relations is nonregular. This in turn implies that all basic relations of $^{\Phi_{0}}\mathfrak{C}$ have size $|S|/2$.
@@
Acknowledgments
The author MM wishes to thank Villanova University for the hospitality shown to him during the final stages of writing this paper. Additionally, the authors are grateful to R. Lipyanski who informed them of the paper [19], and to Misha Klin for helpful discussions and suggestions. Finally, the authors acknowledge the many thoughtful comments made by the referees which greatly enhanced the quality of our paper.